From the top of tower 100m high, a ball is dropped with an angle of 45^°. If it strikes 100m away from the tower then what will be the displacement of the ball?

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Dear student 
x/100 = Sin 45 
x  displacement 
x = 100 /sqrt(2) 
Displacement = sqrt (10000/2 + 10000) = sqrt (15000) 
Hope this helps 
Good Luck